A Complete Bibliography of Publications in the ICES Journal of Marine Science (1940–1949)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.02

Title word cross-reference

Abundance [Tån49]. Algiers [Ber48a]. along [The49]. année [Gil48]. Arendal [Dan47b]. Artificial [Atk48]. Atlantic [Des49, Jac49, Sme47, Tån49]. Atlantique [Des49]. Autumn [Hel48]. Average [dPN48].

Baltic [Jaä49, Mul48, Ott48]. Bancs [Gil48]. Banks [Gil48]. Bay [Rae48]. Belges [Gil48]. Belgian [Gil48]. Bibliography [Ano47a, Ano48a, Ano48b, Ano49a]. Biological [Atk48]. branchialis [Cap47]. Breeding [Tån49]. Buen [Rod48].

C [Tes48]. Côte [Des49]. Callionymus [Let49a]. Caractères [Let49a]. Card [Sme48]. Card-Index [Sme48]. Carl [Hjo48a]. carried [Vib48]. Cas [Let49b]. Cases [Let49b]. certain [Fon48]. certaines [Fon48]. Chalutiers [Gil48]. Channel [Sme47]. Characteristic [Jac49]. Characteristics [Let49a]. chaudes [Ber48b]. chez [Des49, Let49b, dPN48]. cimbrius [Man48]. Clupea [Man48]. Clupéidés [dPN48]. Clupeids [dPN48]. Coast [Des49, Hel48, Dan47a, Dan48b]. Coccolithus [Ber48b]. Composition [Gil48]. Considerations [Gil48, Gil48]. Continuous [Car47]. Copépode [Cap47]. Cos [Rod48]. Council [Ano47b, Sme48]. cours [Gil48]. Course [The49]. critical [Fon48]. critique [Fon48]. Current [Ano47a, Ano48a, Ano48b, Ano49a, Car47]. cycle [Ber48b].

d’Alger [Ber48a]. dans [Fon48]. D’Arcy [Cla49]. deceased [Ano47b]. Del [Rod48]. Detection [Atk48]. Determination [Ott48, The49]. Dimorphism [Des49]. Dimorphisme [Des49]. d’investigation [Fon48]. Distribution [Man48, Mul48]. diverses [Fon48]. Division [Glo47]. dominant [Ber48b]. dominating [Ber48b]. during [Gil48, Mul48, Par49].

Echometer [Bal48]. Editorial [Ano49b]. effectués [Vib48]. Eggs [Man48]. Español [Gar48]. Establishment [Glo47]. étude [Fon48]. Étudiée [dPN48]. Experiments [Dan48a]. exploited [Gil48]. exploités [Gil48].

Facteurs [Fon48]. factors [Fon48]. Faroe [Rae47]. Faxa [Rae48]. Features [Jac49]. fertilité [Ber48a]. fertility [Ber48a]. Finland [Hel48]. Fish [Bru49, Dan47b, Let49b, Tån49, Fon48]. Fishery [Rae47]. Fishing [Bal48]. Fjord [JK47]. Fladen [Gil48]. Flagella [Ber48b]. Flagellé [Ber48b]. Flødevig [Dan47b]. Flødevigen [Dan48a]. Food [Bru49]. fragilis [Ber48b]. Francaise [Des49]. France [Vib48]. Free [Atk48]. French [Ber48b, Ber48a, Des49, Fon48, Gil48, Let49a, Let49b, dPN48, Vib48].

Gadides [Cap47]. Gadus [Man48]. Gdansk [Man48]. Gilson [Koc48]. Growth [Dan47a, Dan48b]. Growth-Rate [Dan47a]. Guide [Ott48]. Gulf [Man48]. Gullmar [JK47]. Gustave [Koc48]. Gut [Gil48].

H. [Tes48]. Haddock [Par49]. Harengs [Gil48]. Hatchery [Dan47b, Dan48a]. Heinrici [Hjo48a]. Herring [Bal48, Dan47a, Gil48]. Hjort [And49]. Hydrographic [Glo47]. Hydrographical [Sme48].

Iberian [dPN48]. Ibérique [dPN48]. Iceland [Rae48]. Index [Ano48c, Sme48]. Individual [Hjo48b]. Inflow [Sme47]. Institute [Gar48]. Instituto [Gar48]. Instruments [Fra48]. internal [Fon48]. internes [Fon48]. investigation [Fon48].

Jacob [Knu48]. Jacobsen [Knu48]. Johan [Gra48, And49]. joué [Fon48].

L. [Cap47, Hel48, Man48]. large [Ber48a]. Larvae [Man48]. Larval [Bru49]. Layer [Car47]. Lemon [Rae47, Rae48]. Lernaeocera [Cap47]. Lohm [Ber48b]. lyra [Let49a].

Mackerel [Dan48b, Hel48]. Manipulating [Fra48]. marine [Ber48a]. Marking [Jaä49]. Matters [Car47]. Measuring [Car47]. Members [Ano47b]. mers [Ber48b]. méthodes [Fon48]. Methods [The49, Fon48]. Michal [Rus48]. Middle [Mul48]. migrations [Fon48]. morrhua [Man48]. Moyenne [dPN48].

Navy [Glo47]. Nellemose [Mau48]. Nematodes [Bru49]. North [Dan47a, Jac49, Par49, Sme47, Tån49]. Norway [Dan47b]. Norwegian [Dan48b]. Note [Atk48, Let49a, Sme48, Let49a]. Notes [Ano49b].

Observations [Car47]. Oceanografia [Gar48]. Oceanography [Glo47, Gar48]. Odon [Rod48]. off [Ber48a]. Office [Glo47]. Onos [Man48]. Organisms [Fra48]. Orkney [Sme47]. Oscar [Rol48]. other [Car47]. Otto [Tho48]. Ove [Nie49].

Parasite [Cap47]. Paulsen [Nie49]. Peter [Knu48]. Pettersson [Tho48]. Places [Tån49]. Plankton [Fra48]. played [Fon48]. Poissons [Let49b, Fon48]. Potassium [Atk48]. Potassium-Free [Atk48]. Practical [Car47]. préliminaires [Ber48a]. Preliminary [Ber48a]. Programme [Car47]. Proposals [Car47].

Quantitative [Man48]. Quelques [Gil48, Let49b, Des49].

rôle [Fon48]. Racial [dPN48]. Raciale [dPN48]. Rapid [The49]. Rate [Dan47a]. Rearing [Dan48a]. Recherches [Ber48b, Ber48a]. Records [JK47]. Red [Tån49]. Redeke [Tes48]. Reference [Rae47, Rae48]. Region [Jaä49]. Related [Car47]. Relevant [Car47]. Remarks [Car47]. Renaissance [Hjo48b]. research [Ber48a, Ber48b]. Results [Jaä49]. Review [Jaä49, Rae47]. Reviews [Ano48d, B.49, Bar49, Bur48, F.48a, F.48b, F.49, Fag49, Gar47, Gro47, H.47a, H.47b, H.48a, H.48b, H.48c, Luc48, Mar48, McE48, OC47, Ort47, Pet47, R.47a, R.47c, R.47b]. role [Fon48]. Rolf [Gra48].

Salmon [Jaä49, Ott48, Vib48]. Sardine [dPN48]. Saumon [Vib48]. Scomber [Hel48]. scombrus [Hel48]. Scottish [Rae47]. Sea [Atk48, Dan47b, Par49, Sme47]. Seafish [Dan48a]. seas [Ber48b]. Sebastes [Tån49]. Secondaires [Let49a]. Secondary [Let49a]. Sexual [Let49a, Des49]. sexuel [Des49]. Sexuels [Let49a]. Shetland [Sme47]. Ship [The49]. Siedlecki [Rus48]. Significance [dPN48]. Signification [dPN48]. Simple [Fra48]. since [Ano47b]. Skagerak [Dan48b]. Small [Fra48]. Sole [Rae47]. Soles [Rae48]. Some [Fra48, Jac49, Des49, Gil48, Let49b]. South [Hel48]. Spanish [Gar48, Gar48]. Spawning [Dan48b]. Special [Rae47, Rae48]. Spectroscopic [Atk48]. Sprat [Mul48]. sprattus [Man48]. States [Glo47]. Stocks [Par49]. Studied [dPN48]. study [Fon48]. Summer [Mul48]. Sund [Rol48]. Surface [Jac49].

Téléostéens [Des49]. Teleosts [Des49]. Temperature [Jac49, The49]. Tératologiques [Let49b]. Terological [Let49b]. Théodore [lD48]. Thermograms [The49]. Thick [Car47]. Thick-Layer [Car47]. Thompson [Cla49]. Tissier [lD48]. Transparency [JK47]. Travaux [Vib48]. Trawlers [Gil48]. Trout [Ott48].

United [Glo47]. Use [Bru49].

Variabilité [dPN48]. Variability [dPN48]. Variation [Jac49]. various [Fon48]. Vertebral [dPN48]. Vertébrale [dPN48]. Volume [Ano48c].

Wales [Dan47a]. warm [Ber48b]. Water [Atk48, Sme47]. Weathers [Car47]. Wentworth [Cla49]. Whitebait [Dan47a]. Wilhelm [Mau48]. Wind [Car47]. Witting [Gra48]. Work [Vib48].

XV [Ano48c].

year [Gil48]. Years [Rae47]. Young [Dan48b].

References

Andersson:1949:JH

[And49] K. A. Andersson. Johan Hjort 1869–1948. ICES Journal of Marine Science, 16(1):3–8, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

        com/icesjms/article/16/1/3/639577.

Anonymous:1947:CB

[Ano47a] Anonymous. Current bibliography. ICES Journal of Marine Science, 15(1):90–118, July 1947. CODEN ICESEC. ISSN 1054-3139 (print),

        1095-9289 (electronic). URL http://academic.oup.com/icesjms/
        article/15/1/90/553433.

Anonymous:1947:MCD

[Ano47b] Anonymous. Members of the Council deceased since 1939. ICES Journal of Marine Science, 15(1):3–6, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

        oup.com/icesjms/article/15/1/3/553381.

Anonymous:1948:CBa

[Ano48a] Anonymous. Current bibliography. ICES Journal of Marine Science, 15(2):233, April 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/2/233/702998.

Anonymous:1948:CBb

[Ano48b] Anonymous. Current bibliography. ICES Journal of Marine Science, 15(3):351–364, September 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/15/3/351/617594.

Anonymous:1948:IVX

[Ano48c] Anonymous. Index to volume XV. ICES Journal of Marine Science, 15(3):365–367, September 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/15/3/365/617596.

Anonymous:1948:R

[Ano48d] Anonymous. Reviews. ICES Journal of Marine Science, 15(3):345– 347, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/345/617589.

Anonymous:1949:CB

[Ano49a] Anonymous. Current bibliography. ICES Journal of Marine Science, 16(1):119–151, April 1949. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/1/119/639560.

Anonymous:1949:EN

[Ano49b] Anonymous. Editorial notes. ICES Journal of Marine Science, 16(1):16, April 1949. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/1/16/639567.

Atkins:1948:NSB

[Atk48] W. R. G. Atkins. Note on the spectroscopic and biological detection of potassium in sea water and “potassium-free” artificial sea water. ICES Journal of Marine Science, 15(2):169–172, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/15/2/169/702970.

B:1949:R

[B.49] H. B. Reviews. ICES Journal of Marine Science, 16(1):113, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/16/1/
113/639542.

Balls:1948:HFE

[Bal48] Ronald Balls. Herring fishing with the echometer. ICES Journal of Marine Science, 15(2):193–206, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/2/193/702980.

Barnes:1949:R

[Bar49] H. Barnes. Reviews. ICES Journal of Marine Science, 16(1):113–116, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/1/113/639546.

Bernard:1948:RPF

[Ber48a] Francis Bernard. Recherches préliminaires sur la fertilité marine au large d’Alger. (French) [Preliminary research on marine fertility off Algiers]. ICES Journal of Marine Science, 15(3):260–267, September 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/3/
260/617512.

Bernard:1948:RCD

[Ber48b] Francis Bernard. Recherches sur le cycle du Coccolithus fragilis Lohm., Flagellé dominant des mers chaudes. (French) [Research on the cycle of the Coccolithus fragilis Lohm., Flagella dominating warm seas]. ICES Journal of Marine Science, 15(2):177–188, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/2/
177/702975.

Bruun:1949:UNF

[Bru49] Anton Fr. Bruun. The use of nematodes as food for larval fish. ICES Journal of Marine Science, 16(1):96–99, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/16/1/96/639595.

Burgess:1948:R

[Bur48] C. R. Burgess. Reviews. ICES Journal of Marine Science, 15(3):335– 336, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/335/617541.

Capart:1947:LBC

[Cap47] André Capart. Le Lernaeocera branchialis (L.), Copépode parasite des gadides. ICES Journal of Marine Science, 15(1):69–76, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/1/
69/553396.

Carruthers:1947:PPC

[Car47] J. N. Carruthers. Practical proposals for a continuous programme of thick-layer current measuring in all weathers with remarks on relevant wind observations and other related matters. ICES Journal of Marine Science, 15(1):13–26, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/1/13/553376.

Clark:1949:DWT

[Cla49] R. S. Clark. D’Arcy Wentworth Thompson 1860–1948. ICES Journal of Marine Science, 16(1):9–13, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/16/1/9/639594.

Daniel:1947:GRW

[Dan47a] D. Sc. R. J. Daniel. The growth-rate of whitebait herring of the North Wales Coast. ICES Journal of Marine Science, 15(1):42–60, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/1/
42/553389.

Dannevig:1947:FSF

[Dan47b] Alf Dannevig. The Flødevig Sea Fish Hatchery, at Arendal, Norway. ICES Journal of Marine Science, 15(1):7–12, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/15/1/7/553400.

Dannevig:1948:REF

[Dan48a] Alf Dannevig. Rearing experiments at the Flødevigen Seafish Hatchery 1943–1946. ICES Journal of Marine Science, 15(3):277–283, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/277/617518.

Dannevig:1948:SGY

[Dan48b] Alf Dannevig. Spawning and growth of young mackerel on the Norwegian Skagerak Coast. ICES Journal of Marine Science, 15(2): 218–220, April 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/2/218/702984.

Desbrosses:1949:DSC

[Des49] Pierre Desbrosses. Dimorphisme sexuel chez quelques téléostéens de la Côte atlantique française. (French) [Sexual dimorphism in some French Atlantic Coast Teleosts]. ICES Journal of Marine Science, 16(1):63–73, April 1949. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/1/63/639587.

Navarro:1948:VSR

[dPN48] Francisco de P. Navarro. La variabilité et la signification raciale de la moyenne vertébrale des Clupéidés Étudiée chez la sardine ibérique. (French) [Variability and racial significance of the vertebral average of Clupeids studied in the Iberian sardine]. ICES Journal of Marine Science, 15(3):318–327, September 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/3/318/617531.

F:1948:Ra

[F.48a] J. H. F. Reviews. ICES Journal of Marine Science, 15(3):336– 337, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/336/617549.

F:1948:Rb

[F.48b] J. H. F. Reviews. ICES Journal of Marine Science, 15(3):337– 338, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/337/617553.

F:1949:R

[F.49] G. P. F. Reviews. ICES Journal of Marine Science, 16(1):116–117, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/1/116/639550.

Fage:1949:R

[Fag49] Louis Fage. Reviews. ICES Journal of Marine Science, 16(1):117–118, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/1/117/639555.

Fontaine:1948:DRJ

[Fon48] Maurice Fontaine. Du rôle joué par les facteurs internes dans certaines migrations de poissons: étude critique de diverses méthodes d’investigation. (French) [The role played by internal factors in certain fish migrations: a critical study of various methods of investigation]. ICES Journal of Marine Science, 15(3):284–294, September 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/3/
284/617523.

Fraser:1948:SSI

[Fra48] J. H. Fraser. Some simple instruments for manipulating small plankton organisms. ICES Journal of Marine Science, 15(2):173–176, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/2/
173/702972.

Gardiner:1947:R

[Gar47] A. C. Gardiner. Reviews. ICES Journal of Marine Science, 15 (1):80–81, July 1947. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/1/80/553407.

Garcia:1948:IEN

[Gar48] Rafael Garcia. Instituto Español de Oceanografia. (Spanish) [The Spanish Institute of Oceanography]. ICES Journal of Marine Science, 15(3):253–259, September 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/15/3/253/617509.

Gilis:1948:QCC

[Gil48] Ch. Gilis. Quelques considérations sur la composition des bancs de harengs exploités par les chalutiers belges sur le Fladen et le Gut au cours de l’année 1946. (French) [Some considerations on the composition of the Herring Banks exploited by the Belgian trawlers on the Fladen and the Gut during the year 1946]. ICES Journal of Marine Science, 15(2):189–192, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/2/189/702977.

Glover:1947:EDO

[Glo47] R. O. Glover. Establishment of a division of oceanography in the United States Navy hydrographic Office. ICES Journal of Marine Science, 15(1):77–78, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/15/1/77/553403.

Granqvist:1948:RJW

[Gra48] Gunnar Granqvist. Rolf Johan Witting. 1879–1944. ICES Journal of Marine Science, 15(2):137–140, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/2/137/702953.

Gross:1947:R

[Gro47] F. Gross. Reviews. ICES Journal of Marine Science, 15(1):86–87, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
15/1/86/553428.

H:1947:Ra

[H.47a] H. W. H. Reviews. ICES Journal of Marine Science, 15(1):82, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/1/
82/553411.

H:1947:Rb

[H.47b] H. W. H. Reviews. ICES Journal of Marine Science, 15(1):82–83, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/1/
82/553414.

H:1948:Ra

[H.48a] H. W. H. Reviews. ICES Journal of Marine Science, 15(3): 338, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/338/617560.

H:1948:Rb

[H.48b] H. W. H. Reviews. ICES Journal of Marine Science, 15(3):338– 341, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/338/617572.

H:1948:Rc

[H.48c] H. W. H. Reviews. ICES Journal of Marine Science, 15(3):341– 342, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/341/617577.

Hela:1948:MSS

[Hel48] Ilmo Hela. Mackerel (Scomber scombrus L.) on the South Coast of Finland in the autumn of 1936. ICES Journal of Marine Science, 15 (2):221–222, April 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/2/221/702987.

Hjort:1948:CH

[Hjo48a] Johan Hjort. Carl Heinrici. 1874–1944. ICES Journal of Marine Science, 15(2):130–131, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/15/2/130/702948.

Hjort:1948:RI

[Hjo48b] Johan Hjort. The Renaissance of the individual. ICES Journal of Marine Science, 15(2):157–168, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/2/157/702967.

Jaarvi:1949:RRS

[Jaä49] T. H. Jaärvi. A review of the results of salmon marking in the Baltic region. ICES Journal of Marine Science, 16(1):100–112, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/16/1/
100/639538.

Jacobsen:1949:SCF

[Jac49] J. P. Jacobsen. Some characteristic features in the variation of surface temperature in the North Atlantic. ICES Journal of Marine Science, 16(1):17–43, April 1949. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/1/17/639572.

Johnson:1947:RTG

[JK47] N. G. Johnson and F. Koczy. Records of transparency in Gullmar Fjord. ICES Journal of Marine Science, 15(1):39–41, July 1947. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/15/1/39/553385.

Knudsen:1948:JPJ

[Knu48] Martin Knudsen. Jacob Peter Jacobsen. 1877–1946. ICES Journal of Marine Science, 15(2):154–156, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/2/154/702964.

Koch:1948:GG

[Koc48] H. J. Koch. Gustave Gilson. 1859–1944. ICES Journal of Marine Science, 15(2):132–134, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/15/2/132/702949.

leDanois:1948:TT

[lD48] Ed. le Danois. Théodore Tissier. 1866–1944. ICES Journal of Marine Science, 15(2):126–129, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/15/2/126/702946.

Letaconnoux:1949:NCS

[Let49a] R. Letaconnoux. Note sur les caractères sexuels secondaires du Callionymus lyra. (French) [Note on secondary sexual characteristics of Callionymus lyra]. ICES Journal of Marine Science, 16(1):59–62, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/1/59/639586.

Letaconnoux:1949:QCT

[Let49b] R. Letaconnoux. Quelques cas tératologiques chez les poissons. (French) [Some terological cases in fish]. ICES Journal of Marine Science, 16(1):50–58, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/16/1/50/639582.

Lucas:1948:R

[Luc48] C. E. Lucas. Reviews. ICES Journal of Marine Science, 15(3):343– 345, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/343/617585.

Mankowski:1948:QDE

[Man48] W. Mankowski. The quantitative distribution of eggs and larvae of Clupea sprattus L., Gadus morrhua L., and Onos cimbrius L. in the Gulf of Gdansk in 1938, 1946, and 1947. ICES Journal of Marine Science, 15(3):268–276, September 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/3/268/617516.

Marshall:1948:R

[Mar48] N. B. Marshall. Reviews. ICES Journal of Marine Science, 15(3):342– 343, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/342/617581.

Maurice:1948:WN

[Mau48] H. G. Maurice. Wilhelm Nellemose. 1890–1944. ICES Journal of Marine Science, 15(2):151–153, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/2/151/702962.

McEwen:1948:R

[McE48] Geo F. McEwen. Reviews. ICES Journal of Marine Science, 15(3): 334–335, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/334/617537.

Mulicki:1948:DSM

[Mul48] Zygmunt Mulicki. The distribution of the sprat in the Middle Baltic during the summer of 1939. ICES Journal of Marine Science, 15 (2):207–217, April 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/2/207/702982.

Nielsen:1949:OP

[Nie49] E. Steemann Nielsen. Ove Paulsen. 1874–1947. ICES Journal of Marine Science, 16(1):14–15, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/16/1/14/639564.

O-C:1947:R

[OC47] J. O.-C. Reviews. ICES Journal of Marine Science, 15(1):87–89, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/1/
87/553430.

Orton:1947:R

[Ort47] J. H. Orton. Reviews. ICES Journal of Marine Science, 15(1):85, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/1/
85/553420.

Otterstrom:1948:BST

[Ott48] C. V. Otterstrøm. Baltic salmon and trout. guide to determination. ICES Journal of Marine Science, 15(2):223–231, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/15/2/223/702991.

Parrish:1949:HSN

[Par49] B. B. Parrish. Haddock stocks in the North Sea during 1946. ICES Journal of Marine Science, 16(1):74–84, April 1949. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/16/1/74/639589.

Pettersson:1947:R

[Pet47] Hans Pettersson. Reviews. ICES Journal of Marine Science, 15(1): 79, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
15/1/79/553406.

R:1947:Ra

[R.47a] E. S. R. Reviews. ICES Journal of Marine Science, 15(1):84, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/1/
84/553416.

R:1947:Rc

[R.47b] E. S. R. Reviews. ICES Journal of Marine Science, 15(1):85–86, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/1/
85/553424.

R:1947:Rb

[R.47c] W. H. R. Reviews. ICES Journal of Marine Science, 15(1):84–85, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/15/1/
84/553418.

Rae:1947:RSL

[Rae47] Bennet B. Rae. Review of the Scottish lemon sole fishery at Faroe with special reference to the years 1933 to 1938. ICES Journal of Marine Science, 15(1):61–68, July 1947. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/1/61/553391.

Rae:1948:LSI

[Rae48] Bennet B. Rae. Lemon soles at Iceland, 1924–1939, with special reference to Faxa Bay. ICES Journal of Marine Science, 15(3):295– 317, September 1948. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/3/295/617528.

Rodriguez:1948:OBC

[Rod48] Rafael Garcia Rodriguez. Odon de Buen y Del Cos. 1863–1945. ICES Journal of Marine Science, 15(2):135–136, April 1948. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/15/2/135/702951.

Rollefsen:1948:OS

[Rol48] G. Rollefsen. Oscar Sund. 1883–1943. ICES Journal of Marine Science, 15(2):147–150, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/15/2/147/702959.

Russell:1948:MS

[Rus48] E. S. Russell. Michal Siedlecki. 1873–1940. ICES Journal of Marine Science, 15(2):144–146, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/15/2/144/702958.

Smed:1947:IAW

[Sme47] Jens Smed. The inflow of Atlantic water into the North Sea through the Orkney–Shetland Channel. ICES Journal of Marine Science, 15 (1):27–38, July 1947. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/15/1/27/553380.

Smed:1948:NCH

[Sme48] Jens Smed. Note on the Council’s hydrographical card-index. ICES Journal of Marine Science, 15(2):232, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/15/2/232/702995.

Taaning:1949:BPA

[Tån49] AÅ. Vedel Tåning. On the breeding places and abundance of the red fish (Sebastes) in the North Atlantic. ICES Journal of Marine Science, 16(1):85–95, April 1949. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/1/85/639591.

Tesch:1948:HCR

[Tes48] J. J. Tesch. H. C. Redeke. 1873–1945. ICES Journal of Marine Science, 15(2):141–143, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/15/2/141/702956.

Theisen:1949:MRT

[The49] E. Theisen. Methods for rapid temperature determination from thermograms along a Ship’s course. ICES Journal of Marine Science, 16 (1):44–49, April 1949. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/1/44/639581.

Thompson:1948:OP

[Tho48] D’Arcy W. Thompson. Otto Pettersson. 1848–1941. ICES Journal of Marine Science, 15(2):121–125, April 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/15/2/121/702945.

Vibert:1948:TEF

[Vib48] R. Vibert. Travaux effectués en France sur le saumon. (French) [Work carried out in France on salmon]. ICES Journal of Marine Science, 15(3):328–333, September 1948. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/15/3/328/617534.